[Ncep.nhc.nco_contacts] TAFB Shift Log -- Technical Edition -- Tue Aug 29, 2017

TAFB Shift Log nhc.tsbadmin at noaa.gov
Tue Aug 29 10:14:50 UTC 2017

TAFB Shift Log for Tue Aug 29 2017 06:14:44 GMT-0400 (EDT)

Shift: Mid-Shift

Pacific Forecaster: J Lewitsky

Atlantic Forecaster: G Rubio

Surface Analyst: C McElroy

Technical Issues Summary:

GFE/NCP issues:

N-AWIPS issues:

Other issues: Security Camera display remains inoperable.

Miscellaneous issues: 0345Z Rainfall product did not run for TS Harvey.  
0345Z Rainfall product for Pot. TC 10 is named "Tropical Storm 10" in the  
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170829/d796c0ca/attachment.html 

More information about the Ncep.nhc.nco_contacts mailing list